A body is moving along a straight line with a constant velocity of 64 km/h in the positive sense.

At t = 0 the position of the body was +1.4 km. Find the position of the body at t = 0.25


1
Expert's answer
2021-10-01T12:22:15-0400

Given:

"v=64\rm \: km/h"

"x_0=1.4\:\rm km"

"t=0.25\:\rm h"


The final position of the body

"x=x_0+vt=1.4+64*0.25=17.4\:\rm km"